# Data Mesh: a Systematic Gray Literature Review

### v1

@@ 3.1. Data as a Product
It should be (relatively) easy to combine or integrate data from different data products into composite data products. The consistent use of open standards and conventions across data products can increase interoperability between data products. Gray literature lists several standards and practices a data mesh architecture should encompass. For example, the open standards for describing data products and their composition, including interfaces, metadata, and (access and management) policies, can ensure the interoperability of data products within and across organizations. In addition, standard metadata fields and business glossary models can guarantee that all data products are described unambiguously. Finally, a common method for data quality modeling can ensure all domains define their products’ data quality concerning the same KPIs (key performance indicators).
ID1.3.3. Natively Accessible.:
All data products should provide a unique, programmatically accessible endpoint. For example, a data product can access its data as a web service or SQL endpoint. The endpoint description can be included in the metadata to enable consumers to connect to the data product automatically.
ID1.3.4. Self-describing.:
Data products should be documented so consumers can use them without further assistance from the product owners or data stewards. Self-descriptions allow the organization to effectively scale out its data sources and consumers as the human dependencies are reduced as much as possible. The gray literature mentions several metadata that make a data product self-describing. For example, interface descriptions can define all access endpoints and the policies that govern data access. The information about product owners and subject matter experts can ensure that the consumers know whom to contact when needing more information related to the data. The schemas for all the data views can be used to describe the data structure. Service level objectives (SLOs) can describe the product’s quality and act as a data contract. Finally, the metadata of data lineage and provenance can show the origin of the data and gives consumers trust in the data product as it becomes clear how the data is created and processed.
ID1.3.5. Understandable.:
The models and the natural language documentation of data products should help potential consumers to understand the products’ semantics, syntactic, and behaviors/usage. For example, a data product description can include the entities in the data the product consumes and produces, the relationship between those entities, and the relationships with the other data products in the mesh.
ID1.3.6. Secure.:
A data product needs to use appropriate measures to ensure data confidentiality, integrity, and availability. The measures include but are not limited to data access control, data encryption at rest and data in transit, and data access audit logs. In the data mesh architecture, while the security policies are primarily managed centrally, their enforcement happens in a decentralized fashion. For example, product owners can authorize consumers to access their products’ interfaces.
ID1.3.7. Trustworthy.:
@@
Data consumers should be able to trust the data coming from a data product. The gray literature provides several guidelines to enhance the trustworthiness of data products. First, product owners can use a methodology provided by the federated governance team to model data quality to define their own product’s quality on several dimensions (e.g., completeness, accuracy, and timeliness). A set of service-level objectives (SLOs) can capture the target quality for each data product and act as a contract between the product owner and the product consumer. Finally, automation can support product developers in ensuring quality throughout the product’s lifetime (e.g., automated monitoring of data quality issues and execution of corrective actions).
ID1.3.8. Valuable.:
A product needs to possess long-term value to the rest of the organization. The value should be continuously monitored and maintained. This can have implications for the deployment and operation of data products. For example, suppose the data product is an ML prediction service. In that case, the performance of the ML model needs to be continuously monitored and updated as necessary to keep its performance acceptable (i.e., MLOps (machine learning operations) (Steidl et al., 2023) should be used).
Table 3. Domain ownership: codes and sources.
Sub Category Atomic Codes
ID Name Frequency ID Name Frequency Unique Sources
2.1 Domain Formation 42 2.1.1 Decentralized Domains 14 S1,S9,S19,S21,S22,S42,S45,S48,S52,S55,S78,S91,S110
2.1.2 Domain-Driven Design Method 28
S2,S11,S36,S37,S39,S45,S46,S53,S55,S58,S71,S75,S78,
S79,S87,S90,S91,S96,S98,S100,S103,S110,S111,S112
2.2
Domain
Responsibilities
59 2.2.1
Data Product
Development and Operation
49
S2,S4,S9,S13,S16,S20-22,S24,S25,S29,S33,S35,S41-43,
S47,S49,S51,S53,S58,S64,S67,S71,S73,S74,S77,S87
2.2.2 Domain Governance 10 S13,S35, S43, S63, S64,S71,S73,S76,S88
3.2. Domain Ownership of Data
Data mesh distributes the ownership of data to the teams or individuals closest to the origin of the data.
ID2.1. Domain Formation.
@@ 3.2. Domain Ownership of Data
Data mesh decomposes the analytical data in an organization into a set of disjoint domains. Autonomous domain teams are responsible for building and managing data products from domain data. In particular, data mesh advocates applying the Domain-Driven Design (DDD) approach to analytical data. DDD decomposes an organization’s business functions and capabilities into domains, enabling the designing of software systems based on models of the underlying domains (Evans and Evans, 2004). A domain model describes selected aspects of a domain using the vocabulary (i.e., abstractions) shared by domain experts, users, and developers. Organizations already use separation of responsibilities by decomposing the organization into different business domains (e.g., products, sales, and marketing). Data products are built around operational and analytical data produced by these business domains. Hence, business domains form logical constituents for the distribution of data ownership and may be decomposed into subdomains to manage their complexity better (see Section 5.2).
ID2.2. Responsibilities.
@@
We identified two key responsibilities for domain teams: building and maintaining data products and carrying out governance activities at the domain level. Domain teams possess in-depth knowledge about domain data, including potential use cases and consumers of data, and the processes and techniques for maintaining the quality of data. Hence, they can determine which domain data assets should be provisioned as data products. They can use the tools and services provided by the self-serve data platform to create, manage and serve data products (see Section 3.4). Domain teams also carry out governance activities to ensure compliance and conformance of data products to relevant regulations, organizational policies, and standards (see Section 3.3).
Table 4. Federated computational governance: codes and sources.
Sub Category Atomic Codes
ID Name Frequency ID Name Frequency Unique Sources
3.1
Global
Governace
74 3.1.1
Define Organization-wide
Standards and Guidelines
30
S1,S3,S4,S9,S15,S18,S19,S21,S25,S33,S37,S38,
S53,S58,S60,S62,S74,S77,S82,S95,S111
3.1.2
Define and Enforce
Global Governance Policies
15 S9,S21,S28,S36,S37,S53,S63,S67,S110,S111
3.1.3
Define Data Quality
Assessment Methodology
9 S1,S3, S25, S35, S45, S47, S58, S64, S111
3.1.4 Business Glossary Modeling 6 S38,S90,S111,S114
3.1.5 Data Mesh Monitoring 7 S1,S35,S57,S76,S81,S110
3.1.6 Creating Incentive Models 7 S1,S25,S88,S98
3.2
Local
Governance
34 3.2.1 Modeling Products’ Data 8 S1,S2,S13,S16,S24,S37,S55
3.2.2 Data Product Access Control 6 S13,S28,S29,S64,S86,S98
3.2.3
Data Product
Compliance and Conformance
6 S28,S64,S74,S86,S111
3.2.4 Data Product Quality Assurance 9 S13,S20,S28,S47,S63,S64,S86,S88,S110
3.2.5 Data Product Monitoring 5 S4,S48,S71,S88,S110
3.3
Automation in
Governance
11 3.3.1 Objectives of Automation 5 S1,S26,S57,S98,S110
3.3.2
Automatable Governance
Action Types
6 S21,S57,S88,S98,S110
3.3. Federated Computational Governance
@@
Data mesh uses a decentralized governance model in which sovereign domains govern their data products. Overarching governance on a global level across the mesh of the data products mainly aims to ensure interoperability between data products and to enforce standardizations and organization-wide policies. Balancing the levels of centralization and decentralization can be challenging in the federated governance model. On the one hand, centralization gives more control to ensure interoperability and compliance. On the other hand, the domains need sufficient autonomy to build data products in line with their business.
ID3.1. Global Governance.
@@
We refer to the governance activities conducted by the federated governance team on the data mesh level as global governance. The federated governance team comprises representatives from the domains, subject matter experts, governance experts, members from the legal department, the platform team, the management, and so on. The common global governance activities include:
ID3.1.1. Defining and Enforcing Organization-wide Standards.:
A key goal of standards is to increase the interoperability of data products. Examples are a shared API description language and a standard language for metadata modeling. Standards are also used to ensure data products adhere to security and privacy requirements modeled by the federated governance team.
ID3.1.2. Defining Governance Policies.:
Data governance policies define rules or guidelines that enforce consistent and correct collection, storage, access, usage, and management of data assets in an organization. For example, the governance policies at the data mesh level can enforce a minimum level of data security throughout the organization or enforce access control based on central identity management.
ID3.1.3. Developing a Methodology to Define and Assess Data Quality.:
The federated governance team is also responsible for creating a methodology to define data quality to be enforced by data product owners. The methodology consists of dimensions contributing to data quality (e.g., timeliness of data and completeness of data). A standardized methodology can ensure that all data products define quality on a common set of quality dimensions with shared definitions.
ID3.1.4. Maintaining a Common Business Glossary.:
@@
Organizations use business glossaries of business terms and their definitions to ensure the same definitions are used throughout the organization when analyzing data. The data mesh principles discourage creating one canonical data model for the whole organization, which goes against decentralization. However, the gray literature recognizes the importance of having a shared definition of concepts used throughout the organization (e.g., a customer identification number) as this contributes to interoperability. The federated governance team is responsible for business glossary modeling.
ID3.1.5. Monitoring Data Mesh.:
The federated governance team should continuously observe and asses the overall health of the data mesh. For this purpose, It can collect the necessary data from the data mesh to calculate the KPIs (Key Performance Indicators), such as levels of interoperability, compliance, and usage of data products.
ID3.1.6. Creating Incentive Models.:
@@
Provisioning data products will add to the domain teams’ normal workload and will likely not fit their traditional job descriptions. For example, a supply chain department, whose main task is to optimize delivery routes, may not be keen on providing high-quality data to other domains. Without proper incentives, domains might not see a reason to build data products. Thus, organizations implementing or transforming to a data mesh architecture should develop an incentive model to motivate domains to provision their data as products and continuously govern the offered products.
ID3.2. Local Governance.
@@
We refer to the governance activities conducted by the product/domain teams on a product/domain level as local governance. The product owner, potentially in combination with other domain members, is responsible for conducting the local governance activities. The gray literature mentions several responsibilities of the local governance team:
ID3.2.1. Managing the Data Models of the Product.:
The product team uses its domain knowledge to define the data model of the product. The model schema can be changed and extended throughout the product lifetime to reflect business capabilities in the domain better.
ID3.2.2. Managing Data Access Control.:
Domain experts have a thorough understanding of the data and can, therefore, best judge who should and should not have access to (which parts of) the data.
ID3.2.3. Managing Compliance and Conformance.:
@@
Like managing access controls, ensuring compliance and conformance throughout the product’s lifetime requires a thorough understanding of the domain data.
ID3.2.4. Managing Data Quality.:
Using the quality assessment methodology defined by the federated governance team, the product team establishes the product’s target quality based on a set of quality dimensions.
ID3.2.5. Monitoring Data Product Health.:
Throughout the product lifetime, the product owner is responsible for monitoring its operational health and ensuring its capacity is sufficient to meet the needs of consumers.
ID3.3. Automation of Governance Activities.
@@
Automation plays a significant role in the federated computational governance model to ensure it can effectively scale throughout an organization. Especially when data leaves the scope of a data product and thus the control of the product owner, automated decisions can be used to enforce policies in downstream consumers. The gray literature gives several examples of actions that can be automated in a data mesh architecture: data classification (e.g., sensitivity labels), data anonymization and encryption, data quality checks, data retention and deletion to ensure compliance, and automated monitoring, including alerts and automatic responses (e.g., scale a resource to meet demand).
Table 5. Self-Serve data platform: codes and sources.
Sub Category Atomic Codes
ID Name Frequency ID Name Frequency Unique Sources
4.1 Objectives 24 4.1.1
Reduce Required
Specialization
8 S1,S4,S9,S25,S43,S72
4.1.2 Increase Efficiency 9 S2,S7,S9,S43,S53,S55,S65,S71,S92
4.1.3
Enable Uniformity
and Interoperability
7 S1,S19,S33,S88,S93
4.2
Building
Platform
19 4.2.1
Central
Platform Team
7 S1,S2,S22,S58,S73,S85,S86
4.2.2 IaC Blueprints 12 S7,S13,S53,S54,S93,S97,S98,S105,S108,S111
4.3
Platform
Components
149 4.3.1 Compute 7 S2,S31,S37,S40,S53,S64
4.3.2 Networking 8 S40,S50,S87,S102,S105
4.3.3 Polyglot Storage 13 S2,S3,S7,S10,S31,S37,S53,S58,S80,S97,S98
4.3.4
Services for Data
Product Components
22 S2,S3,S7,S12,S31,S37,S42,S58,S64,S72,S73,S98,S108
4.3.5 Metadata Repository 8 S31,S47,S64,S71,S83,S90,S113,S114
4.3.6 Product/Data Catalog 15 S2,S4,S12,S31,S32, S35,S37,S64,S71,S80,S81,S88,S98
4.3.7 Distributed Query Engine 6 S16,S31,S32,S37,S98,S110
4.3.8 Monitoring 10 S2,S4,S13,S33,S35,S37,S42,S64,S71,S72
4.3.9 Product Lifecycle Management 10 S2,S4,S21,S31,S53,S64,S111,S113
4.3.10 Security and Privacy 22 S2,S4,S7,S12,S13,S28,S31,S33,S29,S37,S50,S71,S73,S80
4.3.11 Policy Enforcement 14 S1,S2,S4,S21,S26,S32,S53,S57,S82,S88,S98,S110
4.3.12 BI Tools 9 S10,S31,S47,S64,S66,S70,S98,S108
3.4. Self-Serve Data Platform

### v2

@@ 3.1. Principle 1: Data as a Product
It should be (relatively) easy to combine or integrate data from different data products into composite data products. The consistent use of open standards and conventions across data products can increase interoperability between data products. The gray literature mentions several standards and practices a data mesh architecture should encompass (Furia, 2021; Wulveryck, 2021; Riccomini, 2021; Yordanov, 2021; Chinnasamy and Menon, 2021; Dalvik, 2020; Sequeda, 2021; Broda, 2022a, b; Ramos, 2021; Broda, 2022e; Jochen Christ and Harrer, 2022). For example, the open standards for describing data products and their composition, including interfaces, metadata, and (access and management) policies, can ensure the interoperability of data products within and across organizations. In addition, standard metadata fields and business glossary models can guarantee that all data products are described unambiguously. Finally, a common method for data quality modeling can ensure all domains define their products’ data quality concerning the same KPIs (key performance indicators).
Addressable.:
The interfaces of a data product should provide unique, programmatically accessible endpoints. The interfaces include those for ingesting and serving data and discovering and managing data products (Dehghani, 2022a). The gray literature primarily focused on data access endpoints and mentioned examples such as web service APIs (Wulveryck, 2021; Hicks, 2021; Oracle, 2021a; Gugliciello, 2021; Baker, 2021b; Oracle, 2021b; Ramos, 2021; Wolf, 2022; Strengholt, 2022a), storage buckets (Schultze and Wider, 2020b; Nivas Shankar and Hasson, 2021; Vellante, 2021; Oracle, 2021a; Oostra, 2021b; Orr, 2020; Cooke, 2021; Oostra, 2021a; Modi, 2021; Ramos, 2021; AWS, 2021; Velante, 2021), topics in publish-subscribe messaging systems (Schultze and Wider, 2020b; Wulveryck, 2021; Hicks, 2021; Oracle, 2021a; Dehghani, 2019a; Oostra, 2021b; Orwat, 2021; Stirling, 2021; Pungas, 2021; Oostra, 2021a; Baker, 2021b; Raab and Chiapusso, 2021; Oracle, 2021b; Broda, 2022a; Ramos, 2021; Schultze and Wider, 2020a; Jochen Christ and Harrer, 2022), and SQL endpoints (Gillin, 2021; Hicks, 2021; Andrew, 2022).
Natively Accessible.:
Zhamak (Dehghani, 2022a) emphasizes that a data product should provide multiple endpoints to consume its data so consumers can use their preferred methods to access data products. Along the same lines, the gray literature mentioned that the data product needs to support polyglot data output interfaces (Serra, 2021; Oracle, 2021a; Cooke, 2021; Madan, 2021; Oracle, 2021b; Dehghani, 2021a; Strengholt, 2022e).
Understandable.:
The data models and natural language documentation of data products should help potential consumers understand the products’ semantics, syntactic, and behaviors/usage (Dehghani, 2022a). For example, the documentation of a data product can include the entities in the data the product consumes and produces, the relationship between those entities, the relationships with the other data products in the mesh, and the information about access patterns and how to request the access to data (Furia, 2021; Wulveryck, 2021; Stirling, 2021; Lidman, 2021; Adevinta, 2021; Dehghani, 2021a; Ramos, 2022). The information about product owners and subject matter experts can ensure that the consumers know whom to contact when needing more information related to the data. The schemas for all the data views can be used to describe the data structure. Service level objectives (SLOs) can describe the product’s quality and act as a data contract. The metadata of data lineage and provenance can show the origin of the data and give consumers trust in the data product as it becomes clear how the data is created and processed. The literature suggests using common data modeling languages or standards can help improve the understanding of data/messages exchanged between data products (Wulveryck, 2021; Stirling, 2021).
Secure.:
A data product must enforce the necessary security policies to ensure the consumers can use it in a secure and regulatory-compliant way (Dehghani, 2022a). The policies include but are not limited to data access control, data encryption at rest and data in transit, data confidentiality, and data retention. In the data mesh architecture, while the security policies are primarily managed centrally, their enforcement happens in a decentralized fashion (Dehghani, 2022a). The policy-as-code approach, where policies are defined and enforced using code in a high-level language, was recommended by Zhamak (Dehghani, 2022a) and other practitioners (Riccomini, 2021; Oracle, 2021a; Strengholt, 2021; Oracle, 2021b; Jochen Christ and Harrer, 2022; Strengholt, 2022a) for managing security policies of data products.
Trustworthy.:
@@
Data consumers should be able to trust the data coming from a data product. The gray literature provides several guidelines to enhance the trustworthiness of data products. Product owners can use a methodology provided by the federated governance team to model data quality to define their own product’s quality on several dimensions (e.g., completeness, accuracy, and timeliness) (Jarvin Mutatiina and Blaauw, 2020; Nivas Shankar and Hasson, 2021; Moses, 2021; Wolf, 2022). Moreover, the self-serve platform can support product developers in ensuring quality throughout the product’s lifetime (e.g., automated monitoring of data quality issues and execution of corrective actions) (Stirling, 2021; Baker, 2021a). Product owners and consumers can also use a formal contract that defines a set of service-level objectives (SLOs), reflecting the target quality for each data product  (Madan, 2021; Reid, 2021; Ramos, 2022; Strengholt, 2022a). Finally, another approach to improve the trustworthiness of data is to provide data provenance and lineage information to the product consumers so that they can inspect the origin and history of the data they use (Wulveryck, 2021; Cooke, 2021; Thoughtworks, 2021).
Valuable on its own.:
The data offered by a product should possess long-term value to the rest of the organization. Even without being aggregated with other data products, a data product should support meaningful use cases (Dehghani, 2022a). However, the gray literature mostly provides the use cases where the data from one domain becomes valuable to the organization once it is combined or correlated with the data from other domains, for example, user behavior data and sales information (Schultze and Wider, 2020b) and smart device data and enterprise data (Wolf, 2022). The value of a data product should be continuously monitored and maintained. The gray literature mentioned adopting the DataOps methodology (Munappy et al., 2020) for automating the continuous delivery of high-quality data products to consumers (Riccomini, 2021; Limited, 2021; Oracle, 2021b; Ramos, 2021).
3.2. Principle 2: Domain Ownership of Data
A central data team often manages and shares analytical data in traditional data architectures. In contrast, data mesh distributes the ownership of data to the teams or individuals (belonging to the domains in the organizations) closest to the data and knows the data intimately (Dehghani, 2022a). Autonomous domain teams are responsible for building and managing data products from domain data. Zhamak (Dehghani, 2022b, 2019b, a) advocates applying the Domain-Driven Design (DDD) approach (Evans and Evans, 2004) to decompose the analytical data in an organization into data domains that align with organizational units.
Domain Formation.
@@ 3.2. Principle 2: Domain Ownership of Data
The gray literature is in agreement with Zhamak regarding the use of DDD for identifying data domains and their data products (Oracle, 2021a; Rajagopalan, 2021; Orwat, 2021; Engineering, 2021; Stirling, 2021; Lidman, 2021; McKenzie, 2021; Oracle, 2021b; Kumar, 2022; Strenholt and Courtright, 2022; McLarty, 2021; Platter, 2021b; Dehghani, 2021a; Ramos, 2021; Maharana, 2021; Jochen Christ and Harrer, 2022; Strengholt, 2022b). DDD decomposes an organization’s business functions and capabilities into domains, enabling the designing of software systems based on models of the underlying domains (Evans and Evans, 2004). A domain model describes selected aspects of a domain using the vocabulary (i.e., abstractions) shared by domain experts, users, and developers. Organizations already use separation of responsibilities by decomposing the organization into different business domains (e.g., products, sales, and marketing). Data products can be built around operational and analytical data produced by these business domains. Hence, business domains form logical constituents for the distribution of data ownership and may be decomposed into subdomains to manage their complexity better. Section 5.2 discusses the application of DDD for building data products.
Responsibilities.
@@
Zhamak (Dehghani, 2022b, 2019b, a) identified two key responsibilities for domain teams: building and maintaining data products and carrying out governance activities at the domain level. The gray literature also mentions similar responsibilities (Moses, 2020; Furia, 2021; Nivas Shankar and Hasson, 2021; Gillin, 2021; Menning, 2021; Riccomini, 2021; Mitrus, 2021; Hicks, 2021; Thandavarayan, 2021; Oracle, 2021a; Keboola, 2021; Rajagopalan, 2021; Saurabh, 2021; Oostra, 2021b; Snowflake, 2021; Dalvik, 2020; Sequeda, 2021; Orr, 2020; Cooke, 2021; Moses, 2021; Shahrin, 2021; Madan, 2021; Pungas, 2021; Baker, 2021b; Raab and Chiapusso, 2021; Butler, 2021; Thoughtworks, 2020, 2021; Strenholt and Courtright, 2022). Domain teams possess in-depth knowledge about domain data, including potential use cases and consumers of data, as well as the processes and techniques for maintaining data quality. Hence, they can determine which domain data assets should be provisioned as data products. They can use the tools and services the self-serve data platform provides to create, manage, and serve data products (see Section 3.4). Domain teams also carry out governance activities to ensure compliance and conformance of data products to relevant regulations, organizational policies, and standards (see Section 3.3).
3.3. Principle 3: Federated Computational Governance
@@
Data mesh uses a decentralized governance model in which sovereign domains govern their data products (Dehghani, 2022b, 2019b, a). Overarching governance on a global level across the mesh of the data products mainly aims to ensure interoperability between data products and to enforce standardizations and organization-wide policies (Jarvin Mutatiina and Blaauw, 2020; Moses, 2020; Furia, 2021; Wulveryck, 2021; ELEKS, 2021; Lam, 2021; Riccomini, 2021; Thandavarayan, 2021; Keboola, 2021; Limited, 2021; Dehghani, 2019a; Cooke, 2021; Moses, 2021; Bulpin, 2021; one Musa, 2020; Shahrin, 2021; Butler, 2021; Thoughtworks, 2020; Broda, 2022a; McLarty, 2021; Strengholt, 2022c). Balancing the levels of centralization and decentralization can be challenging in the federated governance model. On the one hand, centralization gives more control to ensure interoperability and compliance. On the other hand, the domains need sufficient autonomy to build data products that align with their business.
Global Governance.
@@
We refer to the governance activities conducted by the federated governance team on the data mesh level as global governance. The federated governance team comprises different kinds of representatives such as domain experts, subject matter experts, governance experts, members from the legal department, the platform team, and the management (Dehghani, 2022a). The common global governance activities include:
Defining and Enforcing Organization-wide Standards.:
A key goal of standards is to increase the interoperability of data products (Dehghani, 2022b, 2019b, a). Standards can also used to ensure data products adhere to security and privacy requirements modeled by the federated governance team. Zhamak (Dehghani, 2022a) emphasized the standardization of the interfaces of data products, data and query modeling, and data lineage. The gray literature proposed several standards: W3C data catalog vocabulary for product metadata modeling (Sequeda, 2021), cloud events for event data (Menning, 2021), Open API for product interfaces (Gugliciello, 2021; Jochen Christ and Harrer, 2022), JSON schema for data modeling (Riccomini, 2021; Jochen Christ and Harrer, 2022), and ONNX (Open Neural Network Exchange) for machine leaving models (Jochen Christ and Harrer, 2022).
Defining and Enforcing Data Governance Policies.:
Data governance policies define rules or guidelines that enforce consistent and correct collection, storage, access, usage, and management of data assets in an organization. Zhamak (Dehghani, 2022a) recommends cross-cutting concerns such as regulatory compliance, user identity management, and product ownership management for mesh-level governance policies. Several gray literature articles presented examples of using a central identity and access management (IAM) provider to manage users and their permissions (Nivas Shankar and Hasson, 2021; Maharana, 2021) and a data catalog for automatically discovering and classifying sensitive data (Nivas Shankar and Hasson, 2021; Raab and Chiapusso, 2021; Maharana, 2021).
Developing a Methodology to Define and Assess Data Quality.:
The federated governance team is also responsible for creating a methodology to define data quality to be enforced by data product owners (Dehghani, 2022a). According to literature sources (Jarvin Mutatiina and Blaauw, 2020; Riccomini, 2021; Stirling, 2021; Moses, 2021; Pollock, 2021), a standardized methodology and tools can ensure that all data products define quality on a common set of quality dimensions with shared definitions. A lack of consistent data quality enforcement will lead to data swamps (Rajagopalan, 2021; Orwat, 2021). The practitioners propose using a common format for data quality constraints and results (Thandavarayan, 2021) and metadata rules for enforcing the consistent modeling of concepts shared across domains (Madan, 2021).
Maintaining a Common Business Glossary.:
@@
Organizations use business glossaries of business terms and their definitions to ensure the same definitions are used throughout the organization when analyzing data. The data mesh principles discourage creating one canonical data model for the whole organization, which goes against decentralization (Dehghani, 2022a). However, some gray literature sources recognize the importance of having a shared definition of concepts used throughout the organization (e.g., master and reference data such as customers and currency codes) as this contributes to the interoperability of data products (Dehghani, 2019a; Strengholt, 2020, 2022c, 2022f).
Monitoring Data Mesh.:
The federated governance team should continuously observe and assess the overall health of the data mesh and make necessary interventions. Zhamak (Dehghani, 2022b, 2019b, a) discusses collecting metrics such as lead time to change a data product, the number of data product consumers, and data quality dimensions. They can be used to determine the value of a data product and make interventions such as promoting or demoting products. The gray literature sources also emphasized the importance of defining and measuring KPIs (Key Performance Indicators) such as levels of interoperability, compliance, and usage of data products (Moses, 2020; Equation, 2021; Baker, 2021b; Thoughtworks, 2021; Jochen Christ and Harrer, 2022).
Creating Incentive Models.:
@@
Provisioning data products will add to the domain teams’ normal workload and will likely not fit their traditional job descriptions (Riccomini, 2021). For example, a supply chain department, whose main task is to optimize delivery routes, may not be keen on providing high-quality data to other domains. Without proper incentives, domains might not see a reason to build data products. Thus, organizations implementing or transforming to a data mesh architecture should develop an incentive model to motivate domains to provision their data as products and continuously govern the offered products (Dehghani, 2022b, 2019b, a). Several gray literature articles emphasized the need for aligning the incentives of data producers and data consumers and motivating data producers to generate the operational metadata necessary to assess the quality and trustworthiness of data products (Dehghani, 2022b, 2019b; Nivas Shankar and Hasson, 2021; Gillin, 2021; Hicks, 2021; Limited, 2021; Stirling, 2021).
Local Governance.
@@
We refer to the governance activities conducted by the product/domain teams on a product/domain level as local governance. The product owner, potentially in combination with other domain members, is responsible for conducting the local governance activities. Zhamak (Dehghani, 2022b, 2019b, a) identified several governance activities at the domain level: domain data modeling, data quality assurance, and execution of data access control policies. The gray literature mentions similar responsibilities for the local governance team:
Managing the Data Models of the Product.:
The product team possesses domain knowledge and thus can accurately model the product’s data (Nivas Shankar and Hasson, 2021; Gillin, 2021; Hicks, 2021; Limited, 2021; Stirling, 2021; Madan, 2021). The schemas of such models need to be evolved accordingly to reflect the changes in the domain’s data sources (Nivas Shankar and Hasson, 2021; Baker, 2021a).
Managing Data Access.:
Domain experts thoroughly understand the data and can, therefore, best judge who should and should not have access to (which parts of) the data (Oracle, 2021a; Madan, 2021; Kumar, 2022; Ramos, 2021). In addition, tasks such as collecting data lineage, implementing data retention policies, and encrypting data can also be implemented at the product level (Nivas Shankar and Hasson, 2021; Perlov, 2021; Ramos, 2021).
Managing Compliance and Conformance.:
@@
Like managing access controls, ensuring compliance and conformance throughout the product’s lifetime requires a thorough understanding of the domain data (Dehghani, 2022a; Perlov, 2021; Kumar, 2022). The examples mentioned by the gray literature include the adherence to the data modeling standards (Shahrin, 2021; Butler, 2021) and interoperability standards (Wells, 2021; Pollock, 2021), and ensuring privacy compliance (Perlov, 2021; Ramos, 2021).
Managing Data Quality.:
Using the quality assessment methodology defined by the federated governance team, the product team establishes the product’s target quality based on a set of quality dimensions (Dehghani, 2022a). The data quality tests mentioned in the gray literature include measuring data quality dimensions, validating schemas, and checking for violations of domain-specific business rules (Stirling, 2021; Nguyen, 2021; Shahrin, 2021; Madan, 2021; Thoughtworks, 2021).
Monitoring Data Product Health:
Throughout the lifetime of the product, the product owner monitors its operational health and ensures its capacity to meet consumers’ needs (Moses, 2020; Equation, 2021; Jochen Christ and Harrer, 2022). The gray literature mentions several metrics for assessing the operational health of data products: data quality metrics, uptime, performance, and cost (Moses, 2020; Equation, 2021; Baker, 2021b).
Automation of Governance Activities.
@@
Automation plays a significant role in the federated computational governance model to ensure it can effectively scale throughout an organization (Dehghani, 2022b, 2019b, a). According to gray literature sources (Nguyen, 2021; Ramos, 2021; Jochen Christ and Harrer, 2022), automation is key to coping with the data governance challenges introduced by frequent regulatory changes and increases in product consumers and data sources. They also provided examples for data governance actions that can be automated (Riccomini, 2021; Nguyen, 2021; Thoughtworks, 2021; Ramos, 2021; Jochen Christ and Harrer, 2022): data classification (e.g., sensitivity labels), data anonymization and encryption, data quality checks, data retention and deletion to ensure compliance, and automated monitoring, including alerts and automatic responses (e.g., scale a resource to meet demand).
3.4. Principle 4: Self-Serve Data Platform